OpenStudy (anonymous):

Here is a simple one. Most Calculus texts will have a proof. Suppose \(f'(x) = g'(x)\) for all \(x \in \mathbb{R}\). Prove that \(f(x) - g(x) = c\) with some constant \(c \in \mathbb{R}\) for all \(x\ in \mathbb{R}\)

OpenStudy (mathteacher1729):

\[d/dx ( f- g) = d/dx (c) \] \[f' - g' = 0 \] add g' to both sides \[f' = g' \]

OpenStudy (anonymous):

We know the derivatives are the same, you are trying to prove that this implies that the difference between the functions is a constant function, not the other way around.

OpenStudy (anonymous):

Hint: MVT

OpenStudy (anonymous):

\[f'(x) = \lim_{h\rightarrow0}\frac{f(x+h)-f(x)}{h}, g'(x) = \lim_{h\rightarrow0}\frac{g(x+h)-g(x)}{h}\] \[f'(x)-g'(x) = 0 \iff\] \[\lim_{h\rightarrow0}\frac{[f(x+h)-g(x+h)]-[f(x)-g(x)]}{h} = 0\] would this work at all?

OpenStudy (zarkon):

you are working way too hard

OpenStudy (anonymous):

i know <.< limited knowledge of this is showing unfortunately =/ anywhos, would it work? lol

OpenStudy (anonymous):

\[h(x) = f(x) - g(x)\]\[h'(x) = f'(x) - g'(x) = 0\] Work from here

OpenStudy (anonymous):

You should be using MVT somewhere in the proof.

OpenStudy (anonymous):

If h(x) is not a constant function, how would this contradict the fact that h'(x) = 0, using the mean value theorem.

OpenStudy (anonymous):

If \(h(x)\) is not constant then there is some \(x, y\) such that \(h(x) \neq h(y)\)

OpenStudy (anonymous):

and then... \[\frac{h(x) - h(y)}{x-y}\] wouldnt be equal to 0? im trying to see this lol >.<

OpenStudy (anonymous):

Yep and then that contradicts the fact that \(f'(x) = 0\). There would be some \(c\) such that \(f'(c) \neq 0\)

OpenStudy (anonymous):

right right. i believe i see it now.

OpenStudy (anonymous):

That's all you need to show. Therefore h(x) is constant.

OpenStudy (anonymous):

interesting indeed! thank you for putting up with my sluggishness >.< you too zarkon, thank you.

OpenStudy (anonymous):

MVT is a powerful tool

OpenStudy (anonymous):

ive taken one Real Analysis class so far, and the professor skipped over it. it wasnt like we did reach that chapter or section. He just skipped it completely.

OpenStudy (anonymous):

didnt*

OpenStudy (anonymous):

That's really strange

OpenStudy (anonymous):

he was a horrible professor.

OpenStudy (anonymous):

i hated that class so much. im gonna try and retake it with someone else.

OpenStudy (anonymous):

You can probably cover what you've missed in that class with a good text. I recommend Spivak's text on Calculus.

OpenStudy (anonymous):

oh you posted it lol

OpenStudy (anonymous):

doesnt sound like a bad idea, i'll buy it once my school money comes in :)